At some temperature, the Kc for the reaction 2NOCl(g) ⇌ 2NO(g) + Cl2(g) is 1.6 x 10^-5. Calculate the concentrations of all species at equilibrium if 1.0 moles of pure NOCl is initially placed in a 2.0 L flask
Answer #1

2 NOC (g) = 2 NO (9) + Cl₂ (9) Initially, 1 mol change - 2x +za At equilibrium, (1-2x) mol & mol conc. at equm, 1-24 mol I &= 116*10-5 203 = 1.6 10 1-24 Now - we can write 1-22 ~ 1 203 - 116x10-5 on as - 8x10-6 x = 0.02 So, concentrations of the spe